Qx = 235 - 3Px + 40A - 20U + 8Pw
The number of detailing jobs he gets per month (Qx) depends on the price he charges per job (Px), his monthly advertising expenditures (A) measured in Rs. 1,000s, the regional percentage unemployment rate (U), and the average price charged by local car wash businesses (Pw) for a standard wash and wax.

(d) Derive the demand curve for detailing jobs under current conditions. Calculate the values needed to fill in the demand schedule below and then plot the schedule on the graph provided.

Px: 105 95 85 75 65 55 45 35 25 15

Qx:
1
Expert's answer
2017-12-05T12:48:06-0500
Qx = 235 - 3Px + 40A - 20U + 8Pw
(d) Derive the demand curve for detailing jobs under current conditions. Calculate the values needed to fill in the demand schedule below and then plot the schedule on the graph provided.
A = Rs 65 per detailing job, A = Rs 3.5 per month, U = 7.5%, Pw = Rs 15, then:
Qx = 235 - 3Px + 40*3.5 - 20*0.075 + 8*15 = 493.5 - 3Px.
So, the demand curve for detailing jobs under current conditions is Qx = 493.5 - 3Px.
Px: 105 95 85 75 65 55 45 35 25 15
Qx: 179 209 239 269 299 329 359 389 419 449